Coronado Industries was organized on January 1, 2021. During its first year, the corporation issued 1,600 shares of $50 par value preferred stock and 110,000 shares of $10 par value common stock. At December 31, the company declared the following cash dividends: 2021, $5,400; 2022, $12,800; and 2023, $27,700.Show the allocation of dividends to each class of stock, assuming the preferred stock dividend is 7% and noncumulative. ( bla
Answer #1

Answer-

ALLOCATION OF DIVIDENDS
YEARS 2021 2022 2023
$ $ $
Total dividend 5400 12800 27700
Allocation to preferred stock 5400 5600 5600
Remainder to common stock 0 12800-5600 =7200 27700-5600=22100

Dividends to preferred stock =

(1600 shares*$50 per share)*7%= 5600